In an aqueous solution at 25°C, if [H₃O⁺] = 2.4 × 10⁻⁴ M, then [OH⁻] is:


1
Expert's answer
2022-04-08T12:28:03-0400

Solution:

Kw = [H3O+] × [OH]

The value of Kw in any aqueous solution at 25°C is 1.0×10−14 

Therefore,

[OH] = Kw / [H3O+]

[OH] = (1.0×10−14 ) / (2.4×10−4) = 4.2×10−11

[OH] = 4.2×10−11 M


Answer: [OH] is 4.2×10−11 M